OpenStudy (anonymous):

N - 1/5 = 7/20 Find the value for N? Is the answer 11/20? Im not sure.. thanks

jimthompson5910 (jim_thompson5910):

You are correct, the answer is N = 11/20

OpenStudy (anonymous):

thank u.. its good when i get the answer myself.. thanks for your help

jimthompson5910 (jim_thompson5910):

Assuming the equation is \[\Large N-\frac{1}{5}=\frac{7}{20}\]

jhonyy9 (jhonyy9):

N=(7+4)/20 =11/20

OpenStudy (anonymous):

yes Jim it is

jimthompson5910 (jim_thompson5910):

alright, just checking